# PSDNorm: Test-Time Temporal Normalization for Deep Learning on EEG Signals

### v1

@@ 3.2 Implementation details
4:     $\widehat{{\boldsymbol{\mu}}}^{(j)}\leftarrow$ Mean estimation with eq. (10)
@@
Overall Algorithm
The forward computation of the proposed layer is outlined in Algorithm 1. At train time, the PSDNorm performs three main operations: 1) PSD estimation, 2) running Riemannian barycenter update, and 3) Monge mapping application. At test time, the PSDNorm operates similarly, except it does not update the running barycenter. The PSDNorm is fully differentiable and can be integrated into any deep learning model. Similarly to classical normalization layers, a stop gradient operation is applied to the running barycenter to prevent the backpropagation of the gradient computation through the barycenter.
Filter Size and Computational Complexity
PSDNorm has a unique additional hyperparameter $f$ which is the filter size. It controls the alignment between each feature map and the running barycenter PSD and it is typically chosen in $\{16,32,64\}$. In practice, Fourier transforms are computed using the Fast Fourier Transform (FFT) algorithm. Because of the estimation of PSDs, the complexity of the PSDNorm, both at train and test times, is $\mathcal{O}(Nc\ell f\log(f))$, where $N$ is the batch size, $c$ the number of channels, $\ell$ the signal length, and $f$ the filter size.
@@ 4 Numerical Experiments
In this section, we evaluate the proposed method through a series of experiments designed to highlight its effectiveness and robustness. We first describe the datasets and training setup employed in our study, followed by a performance comparison with existing normalization techniques. Next, we assess the robustness of PSDNorm by focusing on its performance over the hardest subjects in the datasets. Finally, we assess its data efficiency under varying training set sizes. The code will be available on github upon acceptance. The anonymized code is available in the supplementary material.
@@ 4.1 Experimental Setup
To evaluate the effect of normalization layers, we use ten datasets of sleep staging described in Table 1. ABC (Jessie P. et al., 2018), CCSHS (Rosen et al., 2003), CFS (Redline et al., 1995), HPAP (Rosen et al., 2012), MROS (Blackwell et al., 2011), SHHS (Quan et al., 1998), CHAT (Marcus et al., 2013), and SOF (Spira et al., 2008) are publicly available sleep datasets with restricted access from National Sleep Research Resource (NSRR) (Zhang et al., 2018). PHYS (Goldberger et al., 2000) and MASS (O’Reilly et al., 2014) are two other datasets publicly available. Every 30 s epoch is labeled with one of the five sleep stages: Wake, N1, N2, N3, and REM. These datasets are unbalanced in terms of age, sex, number of subjects, and have been recorded with different sensors in different institutions which makes the sleep staging task challenging. We now describe the pre-processing steps and splits of the datasets.
@@
The pre-processing pipeline is widely used in the field (Chambon et al., 2017; Stephansen et al., 2018). The datasets vary in the number and type of available EEG and electrooculogram (EOG) channels. To ensure consistency, we use two bipolar EEG channels, as some datasets lack additional channels. For dataset from NSRR, we select the channels C3-A2 and C4-A1. For signals from Physionet and MASS, we use the only available channels Fpz-Cz and Pz-Oz. The EEG signals are low-pass filtered with a 30 Hz cutoff frequency and resampled to 100 Hz. All data extraction and pre-processing steps are implemented using MNE-BIDS (Appelhoff et al., 2019) and MNE-Python (Gramfort et al., 2013).
Table 1: Characteristics of the datasets used in the experiments. The 10 sleep-staging datasets are divided into 7 training datasets, where some subjects are left out for validation and testing, and 3 test datasets that remain entirely unseen during training. Each training dataset is further split into training (64% of subjects), validation (16%), and test (20%) sets.
Dataset Subjects Records Age $\pm$ std Sex % (F/M)
Train ABC 44 117 48.8 $\pm$ 9.8 43/57
CCSHS 231 231 17.7 $\pm$ 0.4 50/50
CFS 681 682 41.7 $\pm$ 20.0 55/45
HPAP 147 147 46.5 $\pm$ 11.9 43/57
MROS 1631 1631 76.4 $\pm$ 5.5 0/100
PHYS 70 132 58.8 $\pm$ 22.0 33/67
SHHS 792 919 63.1 $\pm$ 11.2 52/48
Test MASS 61 61 42.5 $\pm$ 18.9 55/45
CHAT 1230 1635 6.6 $\pm$ 1.4 52/48
SOF 434 434 82.8 $\pm$ 3.1 100/0
Dataset Split
The datasets in Table 1 are divided into train (upper part) and test (lower part) groups. The training datasets are used for both training and testing (with left-out subjects), while the test datasets serve only for testing, i.e., are unseen during training. This split allows us to evaluate the model’s generalization across diverse age groups, sex distributions, and sensors placements. CHAT focuses on children, SOF on elderly women, and MASS on adults. Training datasets are divided into training, validation, and testing sets with a 64%/16%/20% subject split. All experiments are conducted using this split.
@@
Several architectures have been proposed in the literature for sleep staging (Chambon et al., 2017; Stephansen et al., 2018; Phan et al., 2022, 2023). In this work, we use the U-Sleep architecture (Perslev et al., 2021), a state-of-the-art variant of the U-Time model designed for large-scale training and robustness across datasets. The used implementation is from the Braindecode library (Schirrmeister et al., 2017). We use the Adam optimizer (Kingma, 2014), with a learning rate of $10^{-3}$, to minimize the cross-entropy loss. The batch size is set to 64 and the early stopping is done on the validation set with a patience of 5 epochs. The inputs are sequences of 17’30s with a stride of 5’ of the full night recording.
@@
During testing, the model processes sequences of 17’30s with a stride of 5’. Performance is evaluated using the weighted F1 score, computed on the central 7’30s prediction of the sequence to minimize edge effects.
@@
We compare the proposed PSDNorm with three normalization strategies: TMA pre-processing, BatchNorm, and InstanceNorm. TMA is applied as a pre-processing step while retaining BatchNorm in the architecture. Alternatively, BatchNorm layers are replaced in the first three convolutional layers with either PyTorch’s InstanceNorm (Paszke et al., 2019) (using default parameters) or PSDNorm. To preserve the receptive field, the filter size of PSDNorm is set to $f$ in the first layer and progressively halved in subsequent layers. Except for the sensitivity analysis, $f$ is set to 16 and $\alpha$ to $10^{-2}$. A detailed analysis of the neural network depth at which PSDNorm is applied is provided in the supplementary material (Section A.1).
@@ 4.2 Numerical results
In this sub-section, we present the results of the experiments to evaluate the effectiveness of PSDNorm. We compare its performance against the other normalization strategies on the test sets of the training datasets and the full test datasets. Then, we assess the robustness of PSDNorm by focusing on its performance over the hardest subjects in the datasets. We also analyze the data efficiency of PSDNorm by examining its performance with different training set sizes. Finally, the sensitivity of the filter size is investigated to understand its impact on the model’s performance.
Table 2: Scores of different methods on the left-out subjects from the training datasets and the full test datasets.
(Left) F1 score across all subjects. (Right) F1@20% score, representing the F1 score for the 20% lowest-performing subjects. The upper section displays results for training datasets with subject left-out i.e., small shift, while the lower section presents results for testing datasets (unseen at training) i.e., large shift. The best scores are highlighted in bold. N/A means too few subjects to compute the F1@20%. The reported standard deviations indicate performance variability across subjects.
F1 Score F1@20% Score
BatchNorm InstanceNorm TMA PSDNorm BatchNorm InstanceNorm TMA PSDNorm
Train ABC 0.79 $\pm$ 0.06 0.79 $\pm$ 0.06 0.70 $\pm$ 0.08 0.78 $\pm$ 0.08 N/A N/A N/A N/A
CCSHS 0.91 $\pm$ 0.04 0.90 $\pm$ 0.05 0.86 $\pm$ 0.07 0.91 $\pm$ 0.04 0.85 $\pm$ 0.03 0.84 $\pm$ 0.02 0.80 $\pm$ 0.09 0.85 $\pm$ 0.03
CFS 0.90 $\pm$ 0.06 0.89 $\pm$ 0.06 0.86 $\pm$ 0.08 0.89 $\pm$ 0.06 0.80 $\pm$ 0.08 0.81 $\pm$ 0.07 0.77 $\pm$ 0.10 0.80 $\pm$ 0.08
HOMEPAP 0.80 $\pm$ 0.08 0.80 $\pm$ 0.08 0.75 $\pm$ 0.09 0.79 $\pm$ 0.08 0.69 $\pm$ 0.05 0.70 $\pm$ 0.05 0.63 $\pm$ 0.10 0.67 $\pm$ 0.06
MROS 0.89 $\pm$ 0.05 0.89 $\pm$ 0.05 0.87 $\pm$ 0.06 0.89 $\pm$ 0.05 0.81 $\pm$ 0.06 0.81 $\pm$ 0.05 0.79 $\pm$ 0.07 0.81 $\pm$ 0.05
PhysioNet 0.84 $\pm$ 0.05 0.82 $\pm$ 0.07 0.69 $\pm$ 0.08 0.83 $\pm$ 0.05 0.75 $\pm$ 0.04 0.70 $\pm$ 0.07 0.67 $\pm$ 0.07 0.75 $\pm$ 0.00
SHHS 0.87 $\pm$ 0.06 0.87 $\pm$ 0.06 0.80 $\pm$ 0.13 0.87 $\pm$ 0.06 0.77 $\pm$ 0.05 0.79 $\pm$ 0.06 0.68 $\pm$ 0.12 0.78 $\pm$ 0.06
Test CHAT 0.75 $\pm$ 0.09 0.76 $\pm$ 0.09 0.74 $\pm$ 0.10 0.79 $\pm$ 0.07 0.61 $\pm$ 0.07 0.64 $\pm$ 0.08 0.69 $\pm$ 0.12 0.70 $\pm$ 0.07
MASS 0.74 $\pm$ 0.11 0.76 $\pm$ 0.10 0.54 $\pm$ 0.14 0.79 $\pm$ 0.08 0.58 $\pm$ 0.07 0.63 $\pm$ 0.11 0.49 $\pm$ 0.13 0.70 $\pm$ 0.07
SOF 0.86 $\pm$ 0.06 0.86 $\pm$ 0.06 0.83 $\pm$ 0.09 0.86 $\pm$ 0.06 0.76 $\pm$ 0.06 0.76 $\pm$ 0.05 0.72 $\pm$ 0.11 0.76 $\pm$ 0.06
Performance Comparison: Training & Testing Datasets
The left part of the Table 2 presents the F1 scores across the different datasets, highlighting the impact of different normalization strategies under varying distribution shifts. For training datasets, where distribution shifts are relatively small, TMA consistently underperforms compared to other normalization layers, with F1 scores ranging from 2% (MROS) to 15% (PhysioNet) lower. This suggests that applying strong pre-processing with a powerful network like U-Sleep is detrimental. In contrast, all normalization layers achieve comparable results, with F1 scores within 1% of each other. On testing datasets, which exhibit larger distribution shifts, PSDNorm outperforms all baselines, particularly on CHAT and MASS, where it improves over the best alternative by at least 3%. Notably, PSDNorm surpasses BatchNorm by 4% and 5% on CHAT and MASS, demonstrating its effectiveness in handling distribution shifts. On SOF, all normalization layers perform similarly, except for TMA, which shows lower performance. These results demonstrate the effectiveness of PSDNorm in handling distribution shifts, emphasizing that incorporating temporal structure into normalization is crucial for achieving robust and generalizable performance across test datasets.
Robustness: Performance over Most Challenging Subjects
Performance variability across subjects is a key challenge in biomedical applications where ensuring consistently high performance—even for the most challenging subjects—is critical. To evaluate model robustness, we use the F1@20% score (Gnassounou et al., 2023), which measures performance on the 20% of subjects with the lowest F1 scores under the BatchNorm baseline. Across all datasets and methods, performance drops significantly for the hardest subjects compared to all subjects, highlighting the difficulty of the task. TMA performs the worst, reinforcing that strong pre-processing can be detrimental when dealing with subject variability. PSDNorm demonstrates a strong improvement over BatchNorm, achieving a 9% and 12% increase in F1@20% on CHAT and MASS, respectively. Compared to InstanceNorm, PSDNorm provides an important increase, with 6% and 7% improvements on CHAT and MASS. These results show the benefit of using temporal structure for normalization, especially for handling the hardest subjects. On SOF, all normalization methods perform similarly, as expected, given that BatchNorm already achieves high performance on this dataset.
To better highlight the robustness of PSDNorm, Figure 3 presents a scatter plot of subject-wise F1 scores comparing BatchNorm and PSDNorm across different target datasets (upper row). PSDNorm improves performance for 87% of subjects in MASS and 78% in CHAT, with the largest gains observed for the hardest subjects, reinforcing its ability to handle challenging cases.
Figure 3: Subject-wise F1 score comparison between BatchNorm and PSDNorm across test datasets and training set sizes. F1 score of each subject with BatchNorm (x-axis) are plotted against those with PSDNorm (y-axis). The top row shows training with 15% of subjects, the bottom with 100%. Dots are blue when the performance is improved with PSDNorm and orange when it is decreased.
Data Efficiency: Effect of Training Set Size
In many real-world scenarios, especially in biomedical applications, labeled data can be limited and expensive to collect, or simply not available for privacy reasons. This experiment assesses the data efficiency of PSDNorm by analyzing how performance scales with the number of training subjects, determining whether it can achieve strong results with fewer samples. To do so, we vary the percentage of subjects used for training across different datasets while maintaining the original dataset distribution. For each percentage, we perform three random splits and report the average F1 score to ensure robustness. Figure 4 illustrates the impact of training set size on the performance of PSDNorm with different $f$ values, as well as BatchNorm. As expected, performance improves with more training subjects. Furthermore, PSDNorm significantly outperforms BatchNorm in small-data settings, especially for larger $f$, where the normalization effect is stronger. At 15% of subjects, PSDNorm achieves an average F1 score of 77% on test datasets, compared to 66% for BatchNorm, demonstrating its superior ability to learn from limited data. Overall, PSDNorm requires four times fewer subjects to reach the same performance as BatchNorm trained on the full dataset. Figure 3 further highlights this improvement, showing that with only 15% of training subjects, PSDNorm enhances performance for 91% of subjects in SOF, 95% in MASS, and 91% in CHAT, confirming its effectiveness in small-data scenarios.
@@
The choice of $f$ in PSDNorm controls the intensity of the normalization: larger $f$ provide stronger normalization, while smaller $f$ allow more flexibility in the model. In Figure 4, we evaluate its impact across different training set sizes and observe a clear trend: when trained on fewer subjects, larger filter sizes yield better performance, whereas smaller filter sizes are more effective with larger datasets. This suggests that with limited data, stronger normalization is beneficial to avoid overfitting, while with more data, a more flexible model is preferred.
Figure 4: Impact of the number of training subjects on the test datasets F1 score. F1 score is computed on the test datasets (i.e., large shift) for BatchNorm (dashed orange line) and PSDNorm with different numbers of components, $f\in\{16,32,64\}$.
5 Conclusion
In this paper, we introduced PSDNorm, a novel normalization layer for deep neural networks that aligns the power spectral density (PSD) of each signal to a geodesic barycenter at every layer where it is applied. Through extensive experiments on multiple sleep staging datasets, we demonstrated that PSDNorm not only enhances performance but also improves robustness and data efficiency, outperforming existing normalization techniques such as BatchNorm, and InstanceNorm.
Beyond time-series, its adaptability makes it a promising candidate for broader tasks, such as audio analysis, images and videos with high variability in the data distribution.
Impact Statement
We believe our work advances Machine Learning by introducing a test-time adaptation method that improves robustness to distribution shifts, particularly in biomedical applications such as EEG analysis. By leveraging temporal context, our approach enables finer subject-specific adjustments, leading to more accurate and reliable predictions. Furthermore, since adaptation occurs at test time without requiring additional training data, our method helps preserve data privacy, making it well-suited for deployment in sensitive biomedical settings.

### v2

@@ 3.2 Implementation details
4:     $\widehat{{\boldsymbol{\mu}}}^{(j)}\leftarrow$ Mean estimation
@@
The forward computation of the proposed layer is outlined in Algorithm 1. At train time, the PSDNorm performs three main operations: 1) PSD estimation, 2) running Riemannian barycenter update, and 3) Monge mapping application. At test time, the PSDNorm operates similarly, except it does not update the running barycenter. The PSDNorm is fully differentiable and can be integrated into any deep learning model. Similarly to classical normalization layers, a stop gradient operation is applied to the running barycenter to prevent the backpropagation of the gradient computation through the barycenter. PSDNorm has a unique additional hyperparameter $F$ which is the filter size. It controls the alignment between each feature map and the running barycenter PSD and it is typically chosen in our experiments between 1 and 17. In practice, the Fourier transforms are efficiently computed using the Fast Fourier Transform (FFT) algorithm. Because of the estimation of PSDs, the complexity of the PSDNorm, both at train and test times, is $\mathcal{O}(Nc\ell F\log(F))$, where $N$ is the batch size, $c$ the number of channels, $\ell$ the signal length, and $F$ the filter size.
PSDNorm as a generalization of InstanceNorm
InstanceNorm applies a per-channel $z$-score over time, subtracting the mean and dividing by the standard deviation—equivalent to whitening under an i.i.d. assumption over time. In contrast, PSDNorm explicitly accounts for temporal structure by estimating the PSD and whitening/re-coloring in the frequency domain. InstanceNorm is recovered as a special case of PSDNorm by setting the filter size to $F=1$ and using the identity as the re-coloring transform instead of the barycentric PSD.
@@ 4 Numerical Experiments
In this section, we evaluate the proposed method through a series of experiments designed to highlight its effectiveness and robustness on the clinically relevant task of sleep staging. We first describe the datasets and training setup employed, followed by a performance comparison with existing normalization techniques. Next, we assess the efficiency of PSDNorm by training over varying numbers of subjects per dataset. Finally, we analyze the robustness of PSDNorm against domain shift by focusing on subject-wise performance and different architectures. The code will be available on GitHub upon acceptance. The anonymized code is available in the supplementary material. All numerical experiments were conducted using a total of 1500 GPU hours on NVIDIA H100 GPUs.
@@ 4.1 Experimental Setup
Table 1: Characteristics of the datasets.
Dataset Subj. Rec. Age $\pm$ std Sex (F/M)
ABC 44 117 48.8 $\pm$ 9.8 43%/57%
CCSHS 515 515 17.7 $\pm$ 0.4 50%/50%
CFS 681 681 41.7 $\pm$ 20.0 55%/45%
HPAP 166 166 46.5 $\pm$ 11.9 43%/57%
MROS 2101 2698 76.4 $\pm$ 5.5 0%/100%
PHYS 70 132 58.8 $\pm$ 22.0 33%/67%
SHHS 5730 8271 63.1 $\pm$ 11.2 52%/48%
MASS 61 61 42.5 $\pm$ 18.9 55%/45%
CHAT 1230 1635 6.6 $\pm$ 1.4 52%/48%
SOF 434 434 82.8 $\pm$ 3.1 100%/0%
Total 11032 14710
@@
To evaluate the effect of normalization layers, we use ten datasets of sleep staging described in Table 1. ABC [33], CCSHS [34], CFS [35], HPAP [36], MROS [37], SHHS [4], CHAT [5], and SOF [38] are publicly available sleep datasets with restricted access from National Sleep Research Resource (NSRR) [39]. PHYS [40] and MASS [3] are two other datasets publicly available. Every 30 s epoch is labeled with one of the five sleep stages: Wake, N1, N2, N3, and REM. These datasets are unbalanced in terms of age, sex, number of subjects, and have been recorded with different sensors in different institutions which makes the sleep staging task challenging. We now describe the pre-processing steps and splits of the datasets.
@@
We follow a standard pre-processing pipeline used in the field [41, 42]. The datasets vary in the number and type of available EEG and electrooculogram (EOG) channels. To ensure consistency, we use two bipolar EEG channels, as some datasets lack additional channels. For dataset from NSRR, we select the channels C3-A2 and C4-A1. For signals from Physionet and MASS, we use the only available channels Fpz-Cz and Pz-Oz. The EEG signals are low-pass filtered with a 30 Hz cutoff frequency and resampled to 100 Hz. All data extraction and pre-processing steps are implemented using MNE-BIDS [43] and MNE-Python [44].
Leave-One-Dataset-Out (LODO) Setup and Balancing
We evaluate model performance using a leave-one-dataset-out (LODO) protocol: in each fold, one dataset is held out for testing, and the model is trained on the union of the remaining datasets. From the training data, 80% of subjects are used for training and 20% for validation, which is used for early stopping. The full held-out dataset is used for testing. To assess performance in low-data regimes, we also evaluate a variant in which we subsample at most $N$ subjects per dataset, promoting balanced contributions across training sources. We refer to this configuration as balanced@$N$, with $N$ ranging from 40 to 400. The exact number of subjects per dataset in each case is listed in Appendix Table 3.
Table 2: Balanced Accuracy (BACC) scores on the left-out datasets. The top section reports results in the large-scale setting (using all available subjects), while the bottom section presents results in the medium-scale setting (balanced@400). For each row, the best score is highlighted in bold, and standard deviations reflect training variability across 3 random seeds.
BatchNorm LayerNorm InstanceNorm PSDNorm(F=5) PSDNorm(F=9) PSDNorm(F=17)
All subjects ABC 78.49 $\pm$ 0.42 77.94 $\pm$ 0.31 78.83 $\pm$ 0.59 78.56 $\pm$ 0.67 78.73 $\pm$ 0.32 78.60 $\pm$ 0.28
CCSHS 88.79 $\pm$ 0.21 87.51 $\pm$ 0.77 88.75 $\pm$ 0.04 88.56 $\pm$ 0.36 88.48 $\pm$ 0.07 88.52 $\pm$ 0.19
CFS 84.97 $\pm$ 0.37 84.29 $\pm$ 0.67 85.73 $\pm$ 0.29 85.42 $\pm$ 0.09 85.25 $\pm$ 0.19 85.28 $\pm$ 0.12
CHAT 64.72 $\pm$ 3.94 64.36 $\pm$ 0.40 68.86 $\pm$ 2.49 70.57 $\pm$ 1.24 70.36 $\pm$ 2.22 70.71 $\pm$ 2.15
HOMEPAP 76.39 $\pm$ 0.29 75.23 $\pm$ 0.78 76.70 $\pm$ 0.35 76.72 $\pm$ 0.27 76.93 $\pm$ 0.10 77.02 $\pm$ 0.32
MASS 73.71 $\pm$ 0.62 71.39 $\pm$ 3.00 72.12 $\pm$ 0.70 72.51 $\pm$ 1.68 71.34 $\pm$ 2.68 71.61 $\pm$ 0.71
MROS 81.30 $\pm$ 0.25 80.44 $\pm$ 0.29 81.49 $\pm$ 0.18 81.57 $\pm$ 0.34 81.35 $\pm$ 0.17 81.30 $\pm$ 0.13
PhysioNet 76.13 $\pm$ 0.57 75.12 $\pm$ 0.22 76.15 $\pm$ 0.52 75.96 $\pm$ 1.02 76.35 $\pm$ 0.27 76.02 $\pm$ 0.47
SHHS 77.97 $\pm$ 1.46 75.98 $\pm$ 0.48 79.05 $\pm$ 0.89 79.14 $\pm$ 1.01 79.33 $\pm$ 0.62 79.12 $\pm$ 0.11
SOF 81.33 $\pm$ 0.54 81.82 $\pm$ 0.79 81.98 $\pm$ 0.22 82.50 $\pm$ 0.34 82.15 $\pm$ 1.00 81.94 $\pm$ 0.65
Mean 78.38 $\pm$ 0.47 77.41 $\pm$ 0.28 78.97 $\pm$ 0.11 79.15 $\pm$ 0.14 79.03 $\pm$ 0.10 79.01 $\pm$ 0.36
Balanced@400 ABC 78.26 $\pm$ 1.33 75.29 $\pm$ 0.81 78.73 $\pm$ 0.42 78.18 $\pm$ 0.68 78.18 $\pm$ 0.91 77.76 $\pm$ 1.00
CCSHS 87.42 $\pm$ 0.16 85.20 $\pm$ 0.48 87.62 $\pm$ 0.42 87.58 $\pm$ 0.30 87.35 $\pm$ 0.52 87.62 $\pm$ 0.48
CFS 84.32 $\pm$ 0.57 81.66 $\pm$ 1.36 84.72 $\pm$ 0.33 84.29 $\pm$ 0.36 84.06 $\pm$ 0.10 84.46 $\pm$ 0.06
CHAT 66.55 $\pm$ 0.88 61.19 $\pm$ 1.16 64.43 $\pm$ 4.41 70.28 $\pm$ 1.70 68.11 $\pm$ 3.94 69.88 $\pm$ 0.46
HOMEPAP 75.25 $\pm$ 0.50 74.86 $\pm$ 0.25 76.47 $\pm$ 0.63 76.83 $\pm$ 0.61 76.61 $\pm$ 0.74 76.49 $\pm$ 0.45
MASS 70.00 $\pm$ 1.91 68.56 $\pm$ 3.33 71.52 $\pm$ 1.13 72.77 $\pm$ 1.09 73.07 $\pm$ 1.30 72.23 $\pm$ 2.40
MROS 80.37 $\pm$ 0.20 78.05 $\pm$ 0.22 80.28 $\pm$ 0.21 80.26 $\pm$ 0.11 80.32 $\pm$ 0.22 80.70 $\pm$ 0.42
PhysioNet 75.81 $\pm$ 0.13 71.82 $\pm$ 2.12 74.68 $\pm$ 0.55 74.82 $\pm$ 2.11 73.77 $\pm$ 1.73 75.09 $\pm$ 0.97
SHHS 76.44 $\pm$ 0.92 75.12 $\pm$ 0.39 78.68 $\pm$ 0.37 78.88 $\pm$ 0.68 77.28 $\pm$ 0.91 78.41 $\pm$ 0.49
SOF 81.08 $\pm$ 1.14 78.70 $\pm$ 0.50 80.68 $\pm$ 1.38 79.49 $\pm$ 0.41 81.44 $\pm$ 0.97 81.07 $\pm$ 0.66
Mean 77.55 $\pm$ 0.34 75.05 $\pm$ 0.28 77.78 $\pm$ 0.46 78.34 $\pm$ 0.42 78.02 $\pm$ 0.67 78.37 $\pm$ 0.38
@@
Sleep staging has inspired a variety of neural architectures, from early CNN-based models [41, 42, 19] to recent attention-based approaches [21, 17, 20]. We evaluate two architectures: U-Sleep [45, 15], a state-of-the-art temporal CNN model designed for robustness and large-scale training, and a newly introduced architecture, CNNTransformer. CNNTransformer combines a lightweight convolutional encoder with a Transformer applied to epoch-level embeddings. It is specifically tailored for two-channel EEG and designed to scale efficiently to large datasets, while remaining minimal in implementation (under 100 lines of code) and training cost (Appendix A). Its design draws inspiration from recent transformer-based models for time series [46], with an emphasis on simplicity and practicality.
We use the Adam optimizer [47] with a learning rate of $10^{-3}$ to minimize the weighted cross-entropy loss, where class weights are computed from the training set distribution. Training is performed with a batch size of 64, and early stopping is applied based on validation loss with a patience of 3 epochs. Each input corresponds to a sequence of 17’30s, with a stride of 10’30s between sequences along the full-night recording.
@@
At test time, the model similarly processes sequences of 17’30s with a stride of 10’30s. Performance is evaluated using the balanced accuracy score (BACC), computed on the central 10’30s of each prediction window. Each experiment is repeated three times with different random seeds, and we report the mean and standard deviation of BACC.
@@
We compare the proposed PSDNorm with three normalization strategies: BatchNorm, LayerNorm, and InstanceNorm. Note that InstanceNorm corresponds to a special case of PSDNorm with $F=1$ and a fixed identity mapping instead of a learned running barycenter. In the following experiments, the BatchNorm layers in the first three convolutional layers are replaced with either PyTorch’s default implementations of LayerNorm, InstanceNorm [48], or PSDNorm. To preserve the receptive field, the filter size $F$ of PSDNorm is used in the first layer and progressively halved in the following ones. We fix the momentum $\alpha$ to $10^{-2}$.
@@ 4.2 Numerical Results
This section presents results from large-scale sleep stage classification experiments. The analysis begins with a comparison of PSDNorm against standard normalization layers—BatchNorm, LayerNorm, and InstanceNorm—on the full datasets. Then, the data efficiency of each method is evaluated under limited training data regimes. Finally, robustness to distribution shift is assessed via subject-wise performance across multiple neural network architectures.
Performance Comparison on Full Datasets
Table 2 (top) reports the LODO BACC of U-Sleep across all datasets, averaged over three random seeds. PSDNorm consistently outperforms all baseline normalization layers—BatchNorm, LayerNorm, and InstanceNorm—achieving the highest mean BACC of 79.15%, which exceeds BatchNorm (78.38%), InstanceNorm (78.97%), and LayerNorm (77.41%) by at least one standard deviation. This performance holds across tested filter sizes, with PSDNorm reaching 79.03% for $F{=}9$ and 79.01% for $F{=}17$. In total, PSDNorm ranks first on 6 out of the 10 datasets. On the challenging CHAT dataset, where all methods struggle, PSDNorm with $F{=}5$ outperforms all other normalizations by more than 2 percentage points, highlighting its robustness under strong distribution shifts. Although InstanceNorm is a strong baseline—outperforming BatchNorm and LayerNorm by at least one standard deviation on average—it is consistently surpassed by PSDNorm in average performance. In contrast, LayerNorm underperforms across the board, achieving the lowest average BACC and never ranking first, confirming its limited suitability for this task.
Efficiency: Performance with 4$\times$ Less Data
Figure 3: Critical Difference (CD) diagram for two architectures on datasets balanced @400. Average ranks across datasets and subjects for USleep and CNNTransformer. Black lines connect methods that are not significantly different.
The PSDNorm layer improves model performance when trained on the full dataset ($\sim$10000 subjects), but such large-scale data availability is not always the case. In many real-world scenarios—such as rare disease studies, pediatric populations, or data collected in constrained clinical settings—labeled recordings are scarce, expensive to annotate, or restricted due to privacy concerns. Evaluating model robustness under these constraints is therefore essential. To this end, we train all models using the balanced@400 setup, which reduces the training data by a factor of 4 compared to the full-data setting. In this lower-data regime, PSDNorm continues to outperform all baseline normalization strategies on 6 out of 10 datasets and achieves higher average BACC across all tested values of $F$. The performance improvement of PSDNorm over the best baseline is more pronounced in this setting: for $F{=}5$, the BACC gain reaches $+0.56\%$, compared to $+0.18\%$ in the full-data setting. Again, the gains exceed one standard deviation. To assess statistical significance, we conducted a critical difference (CD) test [49]. Figure 3 (top) reports the average rank of each method and the corresponding statistical comparisons. The results confirm that PSDNorm significantly outperforms the baselines, underscoring the value of incorporating temporal structure into normalization for robust and data-efficient generalization. The same trend is observed for U-Sleep trained on all subjects (see in Appendix Figure 6). The following experiments focus on the balanced@400 setup.
Robustness Across Architectures
PSDNorm is a plug-and-play normalization layer that can be seamlessly integrated into various neural network architectures. To demonstrate this flexibility, we evaluate its performance on both the U-Sleep and CNNTransformer models. Figure 3 reports the average rank of each normalization method across datasets and subjects for both architectures using datasets balanced@400. In both architectures, PSDNorm with $F{=}5$ achieves the best overall ranking and demonstrates statistically significant improvements over both BatchNorm, LayerNorm and InstanceNorm. Notably, it is also the only method to rank among the top two across both architectures. The results confirm that PSDNorm generalizes well beyond a single architecture and can provide consistent improvements in diverse modeling setups. InstanceNorm performs competitively in some cases but is never significantly better than PSDNorm. The figure also illustrates that PSDNorm with $F{=}17$ does not yield additional benefit in this medium-scale setting, further supporting the conclusion that moderate temporal context is sufficient at this data scale. Detailed numerical scores for CNNTransformer are reported in the supplementary material (Table 7).
@@
Figure 4: Performance of PSDNorm and BatchNorm with varying training set sizes. The BACC score is plotted against the number of training subjects used with U-Sleep.
The choice of $F$ in PSDNorm controls the intensity of the normalization: larger $F$ provide stronger normalization, while smaller $F$ allow more flexibility in the model. In Figure 4, we evaluate its impact across different training set sizes and observe a clear trend: when trained on fewer subjects, larger filter sizes yield better performance (i.e., $F=17$), whereas smaller filter sizes are more effective with larger datasets (i.e., $F={5}$). This suggests that with limited data, stronger normalization helps prevent overfitting, while with more data, a more flexible model is preferred. On average, PSDNorm with $F=5$ offers a good compromise, achieving one of the best performances across all training set sizes.
Performance on the most challenging subjects
Figure 5: Subject-wise BACC comparison on MASS and CHAT (balanced @400). Blue dot means improvement with PSDNorm.
Performance variability across subjects is a key challenge in biomedical applications where ensuring consistently high performance—even for the most challenging subjects—is critical. To highlight the robustness of PSDNorm, Figure 5 presents a scatter plot of subject-wise BACC scores comparing BatchNorm or InstanceNorm vs. PSDNorm across two selected target datasets. CHAT and MASS are two challenging datasets, where the prediction performance is significantly lower than the other datasets. For CHAT, most of the dots are below the diagonal, indicating that PSDNorm improves performance for 91% of subjects against BatchNorm and 99% of subjects against InstanceNorm, with the largest gains observed for the hardest subjects, reinforcing its ability to handle challenging cases. For MASS, PSDNorm improves performance for 75% of subjects against BatchNorm and 69% against InstanceNorm. This demonstrates that PSDNorm is not only effective in improving overall performance but also excels in enhancing the performance of the most challenging subjects.
5 Conclusion, Limitations, and Future Work
This paper introduced PSDNorm, a normalization layer that aligns the power spectral density (PSD) of each signal to a geodesic barycenter. By leveraging temporal correlations, PSDNorm offers a principled alternative to standard normalization layers. Experiments on large-scale sleep staging datasets show that PSDNorm consistently improves performance, robustness, and data efficiency, especially under domain shift and limited-data settings—outperforming BatchNorm, LayerNorm, and InstanceNorm across architectures.
While the results are promising, some limitations remain. PSDNorm introduces a filter size hyperparameter ($F$) that controls normalization strength; although we provide default values that perform well across datasets, selecting it automatically in adaptive settings could be challenging.
Despite these limitations, PSDNorm is flexible and easy to integrate into existing models. Future work includes extending it to other signals such as audio and other biomedical applications.